# A Mathematical Framework for the Problem of Security for Cognition in Neurotechnology

### v1

@@ 2 Projective Holographic Cognition
Projective (or “Quantum”) Cognition typically concerns itself with mental states at the level of complexity accessible to standard experiments in psychology (that is, small numbers of variables for any given context), which is completely reasonable given the historical goals of the mathematical framework.[38, 39, 40, 41, 42] However, in order to make our framework of Cognitive Security more general,666“General” in the sense used by physicists, mathematicians, etc., essentially meaning “capable of describing a wider range of cases”. we must extend it to the higher-dimensional regime. This paper’s first focus will thus be on the definitions of how Projective Cognition extends to high dimensions, and the derivation of some results on its statistical properties both in terms of internal dynamics of computations and final measurements. We draw on the conceptual origins of Hyperdimensional Computing as a model to explain the robustness of both language and of biological neural computations at the level of large [^7] On the order of $10^{3}$ or more. collections of neurons. The same statistical properties theorized to make neural networks so robust to noise will turn out to have consequences for Cognitive Security. We develop this framework in order to facilitate descriptions of the joint effects of the many active cognitive factors which may come into play in a given situation.
@@
    We hypothesize that the statistical properties of cogit holography correspond to the distinguishability of individuals. This will serve as a key source of information in practical applications of Cognitive Security both via the structure of correlations between cogits within an individual person and via patterns in these correlations between individual persons at a group or population level.
Both of these hypotheses are in fact rather tame in that they can be derived from established results. For the first hypothesis, pairwise, triplet-wise, etc. correlations between cogits are known to be effective as a model of interactions among small numbers of cogits – as verified by extensive work in the field of “Quantum”/Projective Cognition. Given this, extension to higher dimensions would still hold in at least the limiting case where one only assesses correlations among small n-tuples101010“n-tuples” being the standard term for a collection of n things. of cogits. We are not asserting that these correlations will extend to tuples of larger n, [^11] Which to our knowledge have not been studied in the literature on Projective Cognition at the scales discussed in this paper. but we would not be at all surprised if this holds true empirically.
@@
2.1 Mathematical Connections to Quantum Computing
In this section we describe an algebra for HDC on many-cogit systems – and thus for many-qubit systems as well. We then show that it offers exceptional robustness even on the qubit scales already achievable by existing quantum computers, and equally on the cogit scales which should be feasible to study in applying cognitive security. The power of HDC comes from both its robustness to noise and its representational richness. While randomly selected points in a hyperspace are not clustered in any way, proportional distances are. This is because of the concentration of probability [^13] I.E. reduction in the distribution’s variance. with the binomial distribution in bit flips between two vectors, and a brief description of the mathematics behind this will be discussed shortly.
A key challenge in quantum computing is the problem of noise in both the measurements and the computations themselves.[49, 50, 51, 52] The field thus has the goal of improving the ratio of physical qubits to logical qubits – the latter being the equivalent number of qubits upon which operations are performed with sufficient robustness. Improvements in error-correcting codes have reduced this requirement significantly, but existing methods still require many physical qubits per logical one, placing significant limitations on what quantum computers are able to accomplish.[53, 51, 54] Many remarkable innovations have been developed, yet the problem remains a significant practical challenge for applied quantum computing.[49, 50, 53, 51]
Robustness is a major problem for a number of other areas of computing as well. Examples range from more algorithmic cases like machine learning to more hardware-driven cases such as ultra low-power hardware and biological computing. While not usually considered in the context of quantum computing, biological forms of computing have a long history of inspiring advancements in both machine learning and computer hardware, as detailed in reviews by MacPherson et al, Ullman, and Hassabis et al, among others.[55, 56, 57] Recent advancements have included the design of hardware meant to mimic many of the core computational processes of neurons, and their efficiency both in terms of speed and power demands has already garnered much attention.[58, 59, 60]
Moving to connect quantum computing with HDC, there are a small number of papers in the area of HDC which reference qubits or quantum concepts, but in each case solely as an analogy and not in a way which is meant to be taken literally. A paper by Bosch et al develops an algorithm for classical hardware which takes the concept of uncertainty before measurement [^14] In quantum mechanics, there is a fundamental unknowability of any physical value prior to it being measured, and values are thus described by distributions. and applies it to machine learning. Prior to being measured, the classical state vectors are floating-point valued, and upon “measurement” they become binary.[61] Another paper by McDonald et al studied the application of HDC to phasors.[62] Yet thus far there has not been any work on applying the principles of HDC – and the blessing of dimensionality – to quantum computing.
We now turn to a brief discussion of the mathematics behind the central advantage of HDC. Using a Hamming distance [^15] That is, the number of bits in a binary vector which do not match those of a second binary vector. as our metric on a binary unit $d$-cube and normalizing distances by $d$, vectors are on average $0.5$ apart from each other. In other words, we are defining distance as a fraction of maximum possible distance: the distance between opposite ends of the space. If one performs the calculation, one finds that for a 10,000 bit space the clustering of distances around 0.5 is so tight that we have less than 1/1,000,000 random vector pairs closer than 0.476 or farther than 0.524. This means that the distances between two random vectors are so rarely close as to be negligible. Thus proximity between vectors can represent substantive information with extremely low probability of obfuscation by unrelated vectors in the space. Additionally, the vectors themselves can have a substantial fraction of their bits flipped and still maintain the original distance properties as a consequence of the blessing of dimensionality. Consider three vectors $\vec{a}$, $\vec{b}$, and $\vec{c}$. In a hyperspace, if $\vec{c}$ starts close to $\vec{a}$ but moves away from it along some set of dimensions, it is with high probability also moving away from $\vec{b}$. The key point is that this property makes hyperdimensional representations extremely robust to noise, and the mathematics will be discussed more thoroughly in. Section 3
Whereas with Localist Representations we must add additional dimensions to cope with additional complexity or a greater number of objects, HDC begins with a fixed, high dimension and one simply adds additional HV for new concepts. It is overwhelmingly likely that these new HV will be quasi-orthogonal. [^16] Meaning that the angle between two vectors is within a small range around $90^{\circ}$. That is, the angle between any two HV will with very high probability be very close to $90^{\circ}$. This means that we do not need to impose any particular restrictions on their placement unless the number of concepts to be encoded by hypervectors is quite large compared with the dimension of the space. The high probability of quasi-orthogonality between random HV is a more precise mathematical statement of the blessing of dimensionality. If we wish to encode a concept which is similar to some others, we may enforce that its distance(s) to one or more other vectors fall within an appropriate region of the space by constructing a random vector whose angles with respect to the others are within some limit of $0^{\circ}$.
HDC has a number of variations, but all of them are fundamentally algebras comprised of HV and certain required operations. [^17] For those unfamiliar with the general theory of algebras: Ordinary algebra and linear algebra are the two most widely known examples of algebras, but many more are possible. Supposing a set of random binary vectors in a $d$-dimensional space, it is easy to show that vectors will, on average, differ by half of their bits and thus have an average hamming distance of $d/2$ from any other random vector. [^18] For each bit, two random vectors have probability $0.5$ of being different. Summing over the dimension of the hypervector, we have $d/2$. However, while the mean remains unchanged, the variance of the distribution becomes increasingly favorable. As we will show in the coming section, this same property holds for hypervectors of qubits – and thus of cogits as well – just as it does for binary HV.
3 Hyperdimensional Computing Algebras
All forms of HDC represent algebras, but there is flexibility in precisely what algebra a given HDC framework will be. Some key requirements are as follows. [^19] A number of review articles discuss this, but here we make use of the very clear summary provided in McDonald et al’s discussion of phasor-based HDC.[63] This is particularly helpful given they construct a complex-valued holography which is helpful for representing hyperdimensional collections of cogits.
3.1 Required Features
Figure 1: Visual summary of the general structure of a Hyperdimensional Computing algebra. Colored boxes correspond to individual variables, with different shades – of orange, green, blue, and purple respectively – denoting different values. The addition, binding, and dynamics operations convert between individual HV. Each darker grey box represents a particular pairing of indices with variables, and permutation rotates these pairings in the manner shown.
3.1.1 Holographic Addition
$s=[a+b+c]$ (1)
Where it must be the case that
$sim(s,a)\sim sim(s,b)\sim sim(s,c)>>sim(s,d)$ (2)
This means that if we know the dictionary/dictionary of symbols, we can construct the set summed over based on which vectors it is most similar to. The sharp clustering of $sim(\cdot,\cdot)$ around $0.5$ for random HV makes this a robust pseudo-inversion (for lack of a better term), rather than a traditional algebraic inverse.
3.1.2 Holographic Binding
$s=$ $[a\otimes b+c\otimes d]$
$a\otimes s=$ $a\otimes[a\otimes b+c\otimes d]$
$=$ $a\otimes a\otimes b+a\otimes c\otimes d$
$=$ $b+\eta$
$=$ $\sim b$
While reversibility of binding is not required for an HDC algebra, self-reversibility or the existence of inverse elements can be found in some versions of them. Since an arbitrary HV combination is nearly always going to be pseudo-orthogonal, whether $\eta$ is the result of a self-reversal or binding with an inverse element, it will for all intents and purposes be random noise compared with our symbol dictionary. Thus, within the context of a symbol dictionary we can robustly invert binding so long as the inverse (or binding operation itself, for a self-inverse) can recover with high probability only one element of a symbol dictionary. That is, recover one region of the hyperspace which corresponds to a specific symbol.
3.1.3 Holographic Permutation
The standard notation for permutation is $\mathrm{\Pi}_{j}$, with $j$ the rotation count. In other words, $\mathrm{\Pi}_{j}$ means that the 00-th index values for all HV have been shifted to the $j$-th index, the $1$st index is now the $j+1$-th, the $d$-th index now the $j-1$-th, and so on. Holographic permutation is thus also invertible, simply by using $\mathrm{\Pi}_{-j}$ or equivalently $\mathrm{\Pi}_{d+1}$.
$(a\otimes b)\otimes(c\otimes d)$ $=(a\otimes d)\otimes(b\otimes c)$
3.2 Complex Holographic Algebra
The following summary of a basic Complex Holographic Algebra is drawn from a paper by McDonald et al, who are to our knowledge the first to develop one in the literature.[62] In a complex holographic algebra the symbol universe is composed of HV not in $\mathbb{R}^{d}$, but $\mathbb{C}^{d}$. From this universe, some collection of regions correspond to different symbols within our dictionary.
3.2.1 Similarity in a Complex Holographic Dictionary
Just as with real-valued HV, cosine similarity is the metric we can use for complex HV. [^21] Recall that by contrast binary HV use Hamming distance. Both will be used for Quantum Holography.
3.2.2 Complex Holographic Addition
The most straightforward version of addition is simple complex number addition, but with a further normalization to retain a unit-magnitude phasor. This results in a loss of some information, making complex HV addition only approximately reversible.
3.2.3 Complex Holographic Binding
Inversion then becomes multiplication with the complex conjugate,
$\psi^{*}\otimes\psi\otimes\varphi=\varphi$ (5)
4 Projective Holography
While the phasor-based HDC algebra developed by McDonald et al. is effective for complex-valued data such as classical analog signals processing, for example,[62] it is insufficient for our purposes. With qubits we must additionally account for whatever property (often spin) is to be mapped to measurement outcomes – and likewise for cogits. The Bloch sphere [^22] A graphical representation of a qubit, where each axis corresponds to one of the three Pauli matrices whose weighted sum (normalized to unit length) describes a qubit. provides a convenient representation for this, and is shown in Figure 2. While describing the statistics of distances between purely random cogits could be quite problematic in some approaches to formulating the problem, the distances are uniformly distributed if assessed in terms of angular distances rather than surface distances. This drastically simplifies the problem of analyzing error-tolerance in Projective Cognition prior to measurement. Measurement error can be assessed similarly to how one would analyze distances between classical HV, but with some additional caveats we will discuss later.
4.1 Cogit Hypervectors
As we construct our formulation of a cogit holography, let us first start with the general form of the core property which makes HV useful: concentration of distance distributions.
On a Bloch sphere (see Figure 2), we define states in terms of azimuthal angles $\theta$ and $\varphi$. For a complete analysis both are usually required, but if our goal is to make statements about distances between states this is not the case. Supposing a set of randomly selected points on the Bloch sphere which are uniformly distributed with respect to angles, we can see that a HV is simply a vector of such states. So a system with many cogits is already capable of implementing a Projective HV, with the usefulness of that representation scaling with the distributional properties we will now discuss.
Figure 2: The Bloch sphere graphically represents qubits – and thus cogits – in terms of one angle $\varphi$ corresponding to the phasor component and a second angle $\theta$ corresponding to relative probabilities of a qubit/cogit being measured as 00 or $1$. The notation |$\ket{0}$ and |$\ket{1}$ are used to represent the states which exist in superposition in a qubit/cogit. Image credit: Wikimedia Commons.[64]
While physically there is a meaningful distinction between $\theta$ and $\varphi$, in terms of abstract distances between states, they may or may not be relevant.
In the most basic case, we could simply take any two phasors pairwise and define the axis perpendicular to the surface arc between them as our momentary axis of interest. Further, we can rotate about this axis such that one phasor’s angle about this axis is zero. It is then clear that the angle between the two will be uniformly distributed on account of those angles being circular, with this new angle $\vartheta$ being uniform across $[0,\pi]$.
We can extend this to all the phasors making up a PHV in much the same way the binomial distribution can be used to study Hamming distance statistics with binary HV. Rescaling angles by $\frac{1}{\pi}$, the sum of the distances can be expressed as an Irwin-Hall distribution with probability distribution function:[65]
$\mathit{f}_{\Theta}(x,n)=\frac{1}{(n-1)!}\sum_{k=0}^{n}(-1)^{k}\binom{n}{k}(n-k)_{+}^{n-1}$ (6)
$(n-k)_{+}$ denotes a function returning 00 if $n-k<0$ and $n-k$ otherwise.
The Gaussian approximation resulting from the blessings of the Central Limit Theorem is $\mathit{f}_{\Theta}(x,n)\sim\mathcal{N}(\mu=\frac{n}{2},\sigma^{2}=\frac{n}{12})$. We then rescale this distribution by $n$ – to normalize over the number of phasors – and $pi$ – to restore the original angle distribution range. Since a linear rescaling will act on the standard deviation rather than the variance, our rescaled values are $\mu=\frac{\pi}{n}\frac{n}{2}=\frac{\pi}{2}$ and $\sigma=\frac{\pi}{n}\frac{\sqrt{n}}{\sqrt{12}}=\frac{\pi}{\sqrt{12n}}$.
$\mathit{f}_{\Theta}(x,n)\sim\mathcal{N}\left(\frac{\pi}{2},\frac{\pi^{2}}{12n}\right)$ (7)
With this of course holding only for $n$ sufficiently large. However, on the scale of hundreds (if not thousands) of phasors, the approximation is quite good.
For a 1,000-dimensional PHV, for example, the normalized angular distances between random PHV fall outside the range of $0.475$ and $0.525$ at a rate of only about $1$ in $10^{6}$. This means that PHV retain the same distance property which makes classical HV so useful. With the normalized standard deviation decreasing as $1/\sqrt{n}$, higher-dimensional PHV result in normalized distances between random PHV which are ever more strongly concentrated around $\pi/2$. Just as HV are robust to machine noise because distance statistics replace the encoding of information in precise state values with encoding via relative distances in a hyperspace, this phasor equivalent provides robustness to noise in quantum gates and aggregate outcomes of cognitive processes themselves.
4.2 Operations for Projective Holographic Algebra
4.2.1 Similarity
We are not aware of any mechanism for assessing the similarity of qubit or cogit HV prior to measurement. However, describing measurements is very straightforward, and also sufficient for us to use the advantageous properties of HDC.
4.2.2 Addition
4.2.3 Binding
Here we can handle phase binding via the mathematical formulation of what are in quantum computing termed Controlled-Phase gates, which will add the phase angles of two cogits. By matching cogits pairwise from two PHV, this extends directly to the many-cogit setting.
4.2.4 Measurement
$Pr[a]=\innerproduct{a}{\psi}$
$Pr[a]=\bra{\psi}M_{a}^{\dagger}M_{a}\ket{\psi}$
4.2.5 Dynamics
|$\ket{\psi^{\prime}}=H\ket{\psi}$
where $H$ is the operator describing the dynamics of the system.
4.2.6 Permutation
@@
5 Statistics of High-dimensional Cogit Vectors
@@
Having discussed a means of linking the mathematics of Hyperdimensional Computing with that of Projective Cognition, we now turn to analyzing the statistical robustness both of dynamics within a system and of measurements of system states. In the context of Projective Cognition, these are – respectively – cognitive processes and assessments of cogit states (via questions, behavioral tests, or any other form of assessment).
5.1 Gate Fidelity Statistics
To assess whether the blessing of dimensionality is preserved in cogits in the presence of noise, we consider as an example the mathematical description of an amplitude-damping gate from quantum computing. In the context of cogits, amplitude-damping is analogous to reducing certainty regarding the probabilities of different possible future states or measurement outcomes. In other words, it damps the cogits along the |$\ket{0}$ |$\ket{1}$ axis on the Bloch sphere. [^23] See: Figure 2
For amplitude-damping gate with Kraus operators [^24] The term for the matrix description of operations which is the one most often used in quantum mechanics and quantum computing. Given only systems on the scale of hundreds of cogits or more are being considered in this paper, to our knowledge there is no reason to opt for the more detailed Lindblad description. This is a technical note which will not be of significance to most readers, but is included for the sake of readers familiar with quantum computing.
$A_{0}$ $=\begin{bmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{bmatrix}$
$A_{1}$ $=\begin{bmatrix}0&\sqrt{\gamma}\\ 0&0\end{bmatrix}$
the expectation of the maximal change of the difference between two states will be along the Bloch sphere z-axis, and will be of magnitude $2\sqrt{\gamma}$. This is because we are considering two cogits whose amplitudes have been damped, and because considering the differences in states along any other axis – whether x, y, or some combination of the three axes – would result in a change in distance of less than $2\sqrt{\gamma}$. Similar calculations can of course be performed for phase or a combination of amplitude and phase. The upper bound on the expectation of distance change is $\delta_{\theta}=\sin^{-1}(2\sqrt{\gamma})$. If – for the sake of cleaner mathematical analysis – we make the simplifying assumption that this noise is approximately equal for all cogits, we can again use the Irwin-Hall distribution, now on the interval $[0,\pi-\sin^{-1}(2\sqrt{\gamma})]$.
Because of how we have framed the distance statistics in terms of geometry on the Bloch sphere, we can directly apply the classical algebra of random variables. As such, the Gaussian approximation of the Irwin-Hall distribution goes from the prior baseline
$f_{\Theta}(x,n)=\mathcal{N}\left(\frac{\pi}{2},\frac{\pi^{2}}{12n}\right)$ (8)
to the post-damping distribution
$f_{\Theta}(x,n)=\mathcal{N}\left(\frac{\pi-\sin^{-1}(2\sqrt{\gamma})}{2},\frac{\pi^{2}}{12n}\right)$ (9)
This then describes the distribution of distances between pairs of PHV which have been randomly distributed across the Bloch sphere and then subjected to amplitude-damping noise. Similar calculations can readily be performed for other noise models such as phase-damping or – for maximal generality – an anisotropic depolarization.
@@ 5.2 Measurement Statistics
We have shown that PHV states possess the appropriate distance statistics, and now move to discuss measurement robustness properties. For a given cogit’s state |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, we have $\text{Pr}(0)=\alpha^{2}$ and $\text{Pr}(1)=\beta^{2}$. For the sake of notational cleanliness in what follows, we will instead use $p=\alpha^{2}$.
@@
From here, we construct a description of the distance probabilities for measurements taken of two random PHV. As with classical hypervector distance statistics, the normalized mean distance would be 0.5. This holds whether the measurement error probabilities are uniform across all cogits – i.e. for $q_{i}=q$ $\forall$ $i=\{1,2,...,n\}$ – or non-uniform, because for random PHV the cogits will be statistically interchangeable. That is, there would be no significance to the pairing of respective measurement error probabilities with cogits because the indexing of the cogits is irrelevant in describing uniformly randomly distributed PHV.
From here, we move to a compound distribution of these mixture distributions, analogous to the move from Bernoulli trial to binomial distribution. In our mixture case, this is described by the function $Pr[k]=\sum_{j=0}^{k}\left[(1-q)p^{j}(1-p)^{1-j}+q(1-p)^{j}p^{1-j}\right]^{n}$ where $j$ is the number of true Bernoulli “successes” (in our context, the number of cogits where the two PHV measurements would have matched in the absence of error), $k$ is the number of measured “successes”, and $n$ is the number of “trials” (dimension of the two PHV).
@@
Applying the central limit theorem to our modified distribution, the probability of two noisy measurements of the same underlying PHV having similarity – that is, $n$ minus the Hamming distance – greater than or equal to some $z$ can be approximated by a Gaussian distribution with the preceding mean and variance. We can approximate the cumulative probability function using the Gaussian error function erf by integrating the Gaussian from $\infty$- ∞ to 00 and then subtracting half the erf function for a given $z\geq 0$. This is equivalent to multiplying the error compliment function, erfc, by $\frac{1}{2}$.
@@
While $p$ was kept in variable form for the sake of the above analysis, for the statistics of uniformly random PHV it will simply be $0.5$ as even an error-free measurement would be equally likely to yield 0 or 1. In other words, we can in some sense integrate out the measurement errors when considering random PHV. Our simplified equation is then
@@
Interestingly, while the measurement errors are very much relevant in the general case, they cancel from the expression when we are considering the distance statistics solely of uniformly random PHV. The $p=0.5$ case in fact turns out to be equal to the statistics for classical hypervector distances in a setting free from error in measurements of the individual Bernoulli trials. For a 1,000-dimensional PHV distances are clustered tightly around $0.5$, and only about $1$ in $10^{6}$ measurements will have a normalized pairwise distance [^26] Recall that in HDC one describes normalized distances as the Hamming distance divided by the dimension of the HV. outside the range $0.425$ and $0.575$.
In fact, the same property holds for assessing fidelity between a true PHV’s probabilities and the measured probabilities from repeated tests. In this case we have the measurements as Bernoulli trials of each cogits, with a success simply a question of whether or not a bit-flip error occurred in the process of measuring. The mean is then $n(1-q)$ and variance $n(1-q)q$, and the distances can be described similarly to the above as
@@
Again using as reference a 1,000-dimensional PHV, with an error probability $q=0.25$ we would find that around $1$ in $10^{6}$ measurements will have a normalized distance from the true measurement that’s less than $0.185$ or more than $0.315$. Yet even at an error rate of $q=\frac{1}{3}$ – with a $1-10^{-6}$ confidence range of $0.26-4.0$ – we can see that a 1,000-dimensional PHV still allows us to distinguish with high statistical confidence between noisy measurements of the same PHV and measurements of a random PHV. The favorable statistical properties of the geometry of classical HV not only hold in terms of distances between PHV measurements, but also in terms of the difference between PHV measurements and ground truth. The former provides foundational utility, and the latter has has the potential to drastically reduce the number of cognitive/psychological tests and/or sequential runs of an algorithm needed. The blessing of dimensionality thus makes it much easier to confidently distinguish between noise vs. substantive differences in the setting of Projective Holographic Computation – and thus Projective Holographic Cognition.
@@
As with the classical binary case, the advantage of HDC in the Projective Holographic Cognition setting scales rapidly with increasing dimensionality. For two random hypervectors, as $n$ increases the probability mass of their normalized Hamming distances increasingly concentrates around $0.5$. So too with PHV. The intuition – if it can be called that – is the for any given random vector of “flipped” bits which moves the measurements away from the probabilities associated with a PHV, the noise will with high probability also move it away from any other random PHV in high-dimensional space. This is the essence of the blessing of dimensionality, and its extension to the case of quantum computing and Projective Holographic Cognition opens the door to potentially radical improvements in the robustness of quantum computation.

### v4

@@ 2.1 Projective Holographic Cognition
Projective (or “Quantum”) Cognition typically concerns itself with mental states at the level of complexity accessible to standard experiments in psychology (that is, small numbers of variables for any given context), which is completely reasonable given the historical goals of the mathematical framework.[Bruza_2015_quantum_cognition_review, Pothos_2022_quantum_cognition_review, Busemeyer_2023_quantum_cognition_models, Yearsley_2016_quantum_cognition_decision_theories, Busemeyer_2015_quantum_cognition_applied_to_psychology] However, in order to make our framework of Cognitive Neurosecurity more general,999“General” in the sense used by physicists, mathematicians, etc., essentially meaning “capable of describing a wider range of cases”. we must extend it to the higher-dimensional regime. This paper’s first focus will thus be on the description of how Projective Cognition extends to high dimensions, and some results on its statistical properties both in terms of internal dynamics of computations and final measurements. We draw on the conceptual origins of Hyperdimensional Computing as a model to explain the robustness of both language and of biological neural computations at the level of large [^10] On the order of $10^{3}$ or more. collections of neurons. The same statistical properties theorized to make neural networks so robust to noise will turn out to have consequences for Cognitive Neurosecurity. We develop this framework in order to facilitate descriptions of the joint effects of the many active cognitive factors which may come into play in a given situation.
@@
    We hypothesize that the statistical properties of cogit holography correspond to the distinguishability of individuals. This will serve as a key source of information in practical applications of Cognitive Neurosecurity both via the structure of correlations between cogits within an individual person and via patterns in these correlations between individual persons at a group or population level.
Both of these hypotheses are in fact rather tame in that they can be derived from established results. For the first hypothesis, pairwise, triplet-wise, etc. correlations between cogits are known to be effective as a model of interactions among small numbers of cogits – as verified by extensive work in the field of “Quantum”/Projective Cognition. Given this, extension to higher dimensions would still hold in at least the limiting case where one only assesses correlations among small n-tuples131313“n-tuples” being the standard term for a collection of n things. of cogits. We are not asserting that these correlations will extend to tuples of larger n, [^14] Which to our knowledge have not been studied in the literature on Projective Cognition at the scales discussed in this paper. but we would not be at all surprised if this holds true empirically.
@@
Figure 1: The Bloch sphere graphically represents qubits – and thus cogits – in terms of one angle $\varphi$ corresponding to the phasor component and a second angle $\theta$ corresponding to relative probabilities of a qubit/cogit being measured as 00 or $1$. The notation |$\ket{0}$ and |$\ket{1}$ are used to represent the states which exist in superposition in a qubit/cogit. Image credit: Wikimedia Commons.[bloch_sphere_wikimedia]
3 Results
3.1 Distance Statistics of Cogit Vectors
In a classical binary HDC, one uses the Hamming distance to describe the similarity of two HV. In other words, taking the absolute difference of the two vectors and summing. Two maximally different HV will have a distance between them equal to the span of the hypercube defining the hyperspece in which those binary HV exist. Taking the same approach for cogits would fail to capture their inherently probabilistic nature, and thus we make use of a more appropriate metric which accounts for features like entanglement and coherence.
A convenient distance metric is the Bures distance, which is derived from fidelity – itself a measure of the similarity of two states. Here we take two cogit HV |$\ket{\psi}$ and |$\ket{\varphi}$ sharing the same basis (i.e. corresponding to the same properties), with respective density matrices $\rho$ and $\sigma$. Fidelity is defined as $\text{Fi}(\rho,\sigma)=|\innerproduct{\psi}{\varphi}|^{2}$ for pure states. Intuitively, this is the simplified case where we assume that we can measure all cogits with any possible relevance. Then the normalized distance metric is yielded by
$\text{b}(\rho,\sigma)=\frac{1}{\sqrt{2}}\text{D}_{\text{Bures}}(\rho,\sigma)=\sqrt{1-\sqrt{\text{Fi}(\rho,\sigma)}}$ (1)
This is a normalized version of the Bures distance, and both it and the derivation of relevant probability distributions are discussed in greater detail in D.1. One finds that this yields a distribution which decreases exponentially as the normalized distance between two cogit HV decreases. For an $N$-cogit state, the distribution is
$\text{Pr}[text{b}(\rho,\sigma)<v]=1-\frac{(2v^{2}-v^{4})^{N-1}}{N-1}$ (2)
For a distance of even $5\%$ less than the maximum possible distance one already sees the kind of robust concentration of probability which is central to hyperdimensional computing. With $N=100$ there is already a $\sim 0.4\%$ chance of even $5\%$ proximity. For a $500$-cogit HV this drops to $\sim 2.9\times 1.7^{-5}$. For a distance of less than $50\%$, one has probabilities of $\sim 2.9\times 10^{-38}$ and $\sim 1.4\times 10^{-182}$ for $100$-cogit and $500$-cogit states respectively.
@@ 3.1 Distance Statistics of Cogit Vectors
Having discussed a means of linking the mathematics of Hyperdimensional Computing with that of Projective Cognition, we now turn to analyzing the statistical robustness of measurements. In the context of Projective Cognition, these are assessments of cogit states (via questions, behavioral tests, or any other form of assessment).
@@ 3.2 Measurement Statistics
We have shown that projective hypervector states possess the appropriate distance statistics, and now move to discuss measurement robustness properties. For a given cogit’s state |$\ket{\psi}=\alpha\ket{0}+\beta\ket{1}$, we have $\text{Pr}(0)=|\alpha|^{2}$ and $\text{Pr}(1)=|\beta|^{2}$. For the sake of notational cleanliness in what follows, we will instead use $p=\alpha^{2}$.
@@
From here, we construct a description of the distance probabilities for measurements taken of two random projective hypervectors. As with classical hypervector distance statistics, the normalized mean distance would be 0.5. This holds whether the measurement error probabilities are uniform across all cogits – i.e. for $q_{i}=q$ $\forall$ $i=\{1,2,...,n\}$ – or non-uniform, because for random projective hypervectors the cogits will be statistically interchangeable. That is, there would be no significance to the pairing of respective measurement error probabilities with cogits because the indexing of the cogits is irrelevant in describing uniformly randomly distributed projective hypervectors.
From here, we move to a compound distribution of these mixture distributions, analogous to the move from Bernoulli trial to binomial distribution. In our mixture case, this is described by the function $Pr[k]=\sum_{j=0}^{k}\left[(1-q)p^{j}(1-p)^{1-j}+q(1-p)^{j}p^{1-j}\right]^{n}$ where $j$ is the number of true Bernoulli “successes” (in our context, the number of cogits where the two projective hypervectors’ measurements would have matched in the absence of error), $k$ is the number of measured “successes”, and $n$ is the number of “trials” (dimension of each projective hypervector).
@@
Applying the central limit theorem to our modified distribution, the probability of two noisy measurements of the same underlying projective hypervector having similarity – that is, $n$ minus the Hamming distance – greater than or equal to some $z$ can be approximated by a Gaussian distribution with the preceding mean and variance. We can approximate the cumulative probability function using the Gaussian error function erf by integrating the Gaussian from $\infty$- ∞ to 00 and then subtracting half the erf function for a given $z\geq 0$. This is equivalent to multiplying the error compliment function, erfc, by $\frac{1}{2}$.
@@
While $p$ was kept in variable form for the sake of the above analysis, for the statistics of uniformly random projective hypervectors it will simply be $0.5$ as even an error-free measurement would be equally likely to yield 0 or 1. In other words, we can in some sense integrate out the measurement errors when considering random projective hypervectors. Our simplified equation is then
@@
Interestingly, while the measurement errors are very much relevant in the general case, they cancel from the expression when we are considering the distance statistics solely of uniformly random projective hypervectors. The $p=0.5$ case in fact turns out to be equal to the statistics for classical hypervector distances in a setting free from error in measurements of the individual Bernoulli trials. For a 1,000-dimensional projective hypervector distances are clustered tightly around $0.5$, and only about $1$ in $10^{6}$ measurements will have a normalized pairwise distance [^17] Recall that in HDC one describes normalized distances as the Hamming distance divided by the dimension of the hypervectors. outside the range $0.425$ and $0.575$.
In fact, the same property holds for assessing fidelity between a true projective hypervector’s probabilities and the measured probabilities from repeated tests. In this case we have the measurements as Bernoulli trials of each cogits, with a success simply a question of whether or not a bit-flip error occurred in the process of measuring. The mean is then $n(1-q)$ and variance $n(1-q)q$, and the distances can be described similarly to the above as
@@
Again using as reference a 1,000-dimensional projective hypervector, with an error probability $q=0.25$ we would find that around $1$ in $10^{6}$ measurements will have a normalized distance from the true measurement that’s less than $0.185$ or more than $0.315$. Yet even at an error rate of $q=\frac{1}{3}$ – with a $1-10^{-6}$ confidence range of $0.26-4.0$ – we can see that a 1,000-dimensional projective hypervector still allows us to distinguish with high statistical confidence between noisy measurements of the same projective hypervector and measurements of a random projective hypervector. The favorable statistical properties of the geometry of classical hypervectors not only hold in terms of distances between projective hypervector measurements, but also in terms of the difference between projective hypervector measurements and ground truth. The former provides foundational utility, and the latter has has the potential to drastically reduce the number of cognitive/psychological tests and/or sequential runs of an algorithm needed. The blessing of dimensionality thus makes it much easier to confidently distinguish between noise vs. substantive differences in the setting of Projective Holographic Computation – and thus Projective Holographic Cognition.
@@
As with the classical binary case, the advantage of HDC in the Projective Holographic Cognition setting scales rapidly with increasing dimensionality. For two random hypervectors, as $n$ increases the probability mass of their normalized Hamming distances increasingly concentrates around $0.5$. So too with projective hypervectors. The intuition – if it can be called that – is the for any given random vector of “flipped” bits which moves the measurements away from the probabilities associated with a projective hypervector, the noise will with high probability also move it away from any other random projective hypervector in high-dimensional space. This is the essence of the blessing of dimensionality, and its extension to the case of quantum computing and Projective Holographic Cognition opens the door to potentially radical improvements in the robustness of quantum computation.